A recipe calls for 2 ingredients and for no more than 15 grams of sugar or 16 grams of salt to be contained by the ingredients. One unit of ingredient A contains 1 gram of sugar, and 1 unit of ingredient B contains 3 grams of sugar. Also, 1 unit of ingredient A contains 4 grams of salt, and 1 unit of ingredient B contains 1 gram of salt. If the cost of 1 unit of ingredient A is $10.00 and the cost of 1 unit of ingredient B is $13.00, what is the maximum amount of money that could be spent on the 2 ingredients?
Answer by stanbon(75887)

A recipe calls for 2 ingredients and for no more than 15 grams of sugar or 16 grams of salt to be contained by the ingredients.
One unit of ingredient A contains 1 gram of sugar,
and 1 unit of ingredient B contains 3 grams of sugar.
Also, 1 unit of ingredient A contains 4 grams of salt,
and 1 unit of ingredient B contains 1 gram of salt.
============================
Sugar Inequality: A + 3B <= 15
Salt Inequality::4A + B <= 16
------
Modify:
A + 3B <= 15
12A+3B <= 48
---
Subtract to get:
11A <=33
A <= 3 (# of ingrediant A's needed)
-----
To solve for "B":
4A + 12B <= 60
4A + B <= 16
----
Subtract and solve for "B":
11B <= 44
B <= 4 (# of ingrediant B's needed)
=============================================
If the cost of 1 unit of ingredient A is $10.00 and the cost of 1 unit of ingredient B is $13.00, what is the maximum amount of money that could be spent on the 2 ingredients?
----
Ans: Max cost for 3A's and 4B's = 3*10 + 4*13 = 30+52 = $82.00
====================
Cheers,
Stan H.
====================